OpenStudy (anonymous):

Please help!!!! Will badge and fan how do i Simplify this expression? 3+i/2-i * 4+i/2+i A: 13-7i/3 B: 11-7i/3 C: 11+7i/5 D: 13+7i/5

OpenStudy (anonymous):

is it \[\frac{3+i}{2-i}\times \frac{4+i}{2+i}\]?

OpenStudy (anonymous):

yeah that is it sorry.

OpenStudy (anonymous):

it is always the case that \((a+bi)(a-bi)=a^2+b^2\) the denominator is \((2-i)(2+i)=2^2+1^2=5\)

OpenStudy (anonymous):

so you can rule out A and B the numerators is \[(3+i)(4+i)\]do you know how to multiply that out?

OpenStudy (anonymous):

11+7i right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok awesome so its C thank you so much that will make this sheet so much easier haha :)

OpenStudy (anonymous):

once again the answer is C it is always C

OpenStudy (anonymous):

yw post if you have any more questions